# Unveiling Group-Specific Distributed Concept Drift: A Fairness Imperative in Federated Learning

### v3

@@ III-C Objective
We present the optimization objective employed to detect and address group-specific distributed concept drift while upholding the privacy of participating clients’ data within the federation. We adopt a multiple global model approach where we aim to discover clusters of clients, each representing a distinct concept, and train each global model accordingly. To achieve this, we incorporate local group-specific loss monitoring on each client. In other words, for each group $s\in S$ we calculate its loss on each global model $\theta_{m}$ in the set of available global models, $GM$, and assign each client to a specific global model, accordingly.
@@
$\begin{split}\min_{\theta_{m}\in GM}f(\theta_{m})^{t}=\frac{\sum_{k\in K}G_{k}(\theta)\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}{\sum_{k\in K}\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}\\ w_{k}^{t}=\arg\min_{m}\sum_{s\in S}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \text{s.t.}\quad\forall\theta_{m}\in GM:\forall k:w_{k}^{t}=m:\forall s\in S:\\ \ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta_{s}\end{split}$ (3)
@@
  • $\theta_{m}\in GM$ refers to a specific global model;
@@
  • $f(\theta_{m})^{t}$ is the weighted sum of local objectives $G_{k}(\theta)$ of clients that identify with global model $\theta_{m}$;
@@
  • $\delta_{s}$ represents a predefined threshold for the acceptable loss difference of group $s$ between two consecutive timestamps on a global model  [^1] As discussed in [14], every client $k$ preserves its entire history of client identities $w_{k}^{t^{\prime}\leq t}$ and local datasets at each timestamp $D_{k}^{t^{\prime}\leq t}$. To mitigate this overhead, clients could opt to retain only a sliding window of the latest timesteps instead.;
Hence, at each timestep, each client identities with the model that yields the lowest sum of group losses, as long as the loss difference of each group $s$ between the current and the previous timestamp does not surpass the threshold for that group, $\delta_{s}$. If there is no such model, then a new global model will be created and added to the set of global models $GM$. Further details will be explained in the next section.
@@ IV FairFedDrift: Fair Federated Learning under Group-Specific Distributed Concept Drift
Our work offers a solution to address the challenge of group-specific distributed concept drift building upon an existing algorithm, FedDrift, presented in [14]. Algorithm 1 presents the pseudo-code of FairFedDrift.  [^2] Source code can be found at: https://github.com/teresalazar13/FairFedDrift.
@@
1:  Initialize global models GM: $\left[\theta_{0}\right]$, number of timesteps: $T$, number of rounds: $R$, number of clients: $K$, client identities $w_{k\in K}^{t=0}=0$, local epochs: $E$, local batch size: $|B|$, local learning rate: $\eta$.
@@
4:        if $\nexists\theta_{m}\in GM:\forall s\in S,\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta$ then
@@
       $\sum_{s\in S}\begin{cases}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \quad\quad\text{if }\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{m})^{D_{k}^{t-1}}_{s}+\delta_{s}\\ \infty\quad\text{else}\end{cases}$
@@
       $\sum_{s\in S}\begin{cases}\ell(\theta_{i})^{\cup_{t^{\prime}}D_{k}^{t^{\prime}}}_{s}-\ell(\theta_{i})^{\cup_{t*}D_{l}^{t*}}_{s}\\ \quad\quad\text{if }\ell(\theta_{i})^{\cup_{t^{\prime}}D_{k}^{t^{\prime}}}_{s}-\ell(\theta_{i})^{\cup_{t*}D_{l}^{t*}}_{s}\leq\delta_{s}\\ \infty\quad\text{else}\\ \end{cases}$
@@
17:        Add new model to $GM$: $\theta_{k}=\frac{\theta_{i}W_{i}+\theta_{j}W_{j}}{W_{i}+W_{j}}$,
@@ IV-1 Multi-model approach
IV-2 Group-specific distributed concept drift detection:
@@ IV-4 Training and Model Averaging
To sum up, the choice of considering the loss of each group $s\in S$ in FairFedDrift, in contrast to FedDrift [14], stems from its effectiveness in handling group-specific concept drift as it inherently better suited to capture the nuances of fairness in these scenarios where specific groups may experience concept drift at different timesteps. This is attributed to the fact that, in instances of group-specific concept drift, a global loss metric may not manifest substantial variations. We elaborate on this aspect in our experiments section, providing a more comprehensive understanding of the dynamics involved.
@@ V Methodology and Experimental Design
V-A Datasets and Distributed Group-Specific Concept Drift Simulation
In our experimental setup, we leverage diverse datasets derived from established benchmarks in machine learning: MNIST [22], FEMNIST [23], and Adult [24]. These datasets were chosen due to their extensive size, making them well-suited for this FL setup, where each client necessitates a substantial number of data instances across different timesteps. In addition, these datasets have also been employed in fairness studies [25, 26], which makes them suitable to our experiments.
To simulate this FL scenario, we partition each dataset into $K$ clients and $T$ timesteps. Furthermore, we introduce group-specific concept drift within the datasets at specific timesteps, enabling the evaluation of the efficacy of our proposed method under evolving data conditions.
@@ V-A1 MNIST-GDrift
The MNIST dataset consists of gray-scale handwritten digit images, commonly used for digit recognition tasks [22]. We introduce one sensitive attribute, $S$, with two groups ($S=1$ and $S=0$), representing distinct image characteristics. For $S=0$ images, we invert the background and digit colors in comparison to standard MNIST images ($S=1$). To simulate two distinctive group-specific concepts (‘B’ and ‘C’) from the original distribution (‘A’), we label-swap images for $S=0$: for drift ’B’, we swap labels ‘0’ and ‘1’; and for drift ‘C’, we swap labels ‘1’ and ‘2’.
@@ V-A2 FEMNIST-GDrift
V-A3 Adult-GDrift
@@
We experiment on the Adult dataset [24] commonly used in fair machine learning research. The prediction task for the Adult dataset involves determining whether a person’s income exceeds \$50K/yr. In this context, race serves as the sensitive attribute, where $S=1$ typically corresponds to White individuals, and $S=0$ represents non-White individuals. To introduce concept drift, we modify features of the samples where $S=0$. For concept ‘A’, the dataset remains unchanged. For concept ‘B’, the target class of non-White individuals who are Male is set to 1 (positive). For concept ‘C’, the target class of non-White individuals whose ‘relationship’ feature is not ‘Husband’ is set to 1.
V-A4 $\alpha$ parameter
@@
The parameter $\alpha$ is a crucial element in our experimental design, determining the size of the unprivileged group relative to the privileged group. The size of the unprivileged group is a result of multiplying the size of the privileged group by $\alpha$. By adjusting the value of $\alpha$, we can control the balance between the groups in our datasets, allowing us to explore different scenarios where one group may be underrepresented or over-represented relative to the other. As such, when $\alpha$ is closer to zero, the relative size of the unprivileged group ($S=0$) is smaller. To generate a specific number of instances depending on the value of $\alpha$ for the MNIST-GDrift and FEMNIST-GDrift datasets, we adjust the number of transformed images accordingly. However, for the Adult-GDrift dataset, the process is more intricate as the original dataset is oversampled, whose default representation is equivalent to $\alpha=0.16$. If $\alpha>0.16$, we generate privileged instances, and if $\alpha<0.16$, we generate unprivileged instances. We employ SMOTE [27] for creating these instances.
@@ V-B Implementation Details
We employ a convolutional neural network (CNN) on the MNIST-GDrift and FEMNIST-GDrift datasets featuring a convolutional layer, a max-pooling layer, a dense layer, and a softmax activation function (multi-class classification). For the Adult-GDrift dataset we employ a fully connected feed-forward neural network with one hidden layer with 10 $tanh$ units, and a sigmoid output neuron (binary classification). We set $K=10$, $T=10$, $\eta=0.1$, and experiment with different values of $\delta_{s}$: $\{0.05,0.1,0.25,0.5,0.75,1.0,1.25,1.5\}$. Furthermore, for the MNIST-GDrift and FEMNIST-GDrift datasets we set $R=10$, $|B|=32$, $E=5$. For the Adult-GDrift dataset we set $R=1$, $|B|=10$, $E=10$.
@@ V-C Metrics
We evaluate the algorithms based on both accuracy (ACC) and fairness (AEQ, OEQ, OPP). Accuracy Equality (AEQ) [28] is a measure of fairness that calculates the ratio of accuracies across two groups ($S=0$ and $S=1$):
@@
Accuracy Equality (AEQ) [28] is a measure of fairness that calculate the ratio of accuracies across different groups:
@@
Overall Equality of Opportunity (OEQ) [29] calculates the ratio of the average true positive rates (TPRs) across all target classes for two groups. Let $P[\hat{Y}=k\mid S=0,Y=k]$ and $P[\hat{Y}=k\mid S=1,Y=k]$ represent the TPRs for group $S=0$ and $S=1$ for the $k^{th}$ target class, respectively. The formula for OEQ is given by:
$\text{OEQ}=\frac{1}{K}\sum_{k=1}^{K}\frac{P[\hat{Y}=k\mid S=0,Y=k]}{P[\hat{Y}=k\mid S=1,Y=k]}$ (5)
where $K$ is the number of target classes.
@@
Overall Predictive Parity (OPP) [30] calculates the ratio of the average positive predictive values (PPVs) across all target classes for two groups. Let $P[Y=k\mid S=0,\hat{Y}=k]$ and $P[Y=k\mid S=1,\hat{Y}=k]$ represent the PPVs for group $S=0$ and $S=1$ for the $k^{th}$ target class, respectively. The formula for OPP is given by:
$\text{OPP}=\frac{1}{K}\sum_{k=1}^{K}\frac{P[Y=k\mid S=0,\hat{Y}=k]}{P[Y=k\mid S=1,\hat{Y}=k]}$ (6)
where $K$ is the number of target classes.
@@
V-D Group-Specific Distributed Concept Drift Scenarios
For our experiments, we design three distinct scenarios of group-specific distributed concept drift, as illustrated in Figure 2. These aim to provide a comprehensive exploration of the temporal and spatial dynamics that can impact fairness and accuracy in FL settings.
V-D1 Baselines
FairFedDrift is compared against three pivotal algorithms: FedAvg, FedDrift [14] and Oracle. The Oracle algorithm serves as an idealized benchmark, equipped with access to concept IDs during training and employing multiple-model training based on ground-truth clustering. Given the unique problem of group-specific distributed concept drift, traditional baselines for fair FL are absent, as explained in the Related Work section. We found that these achieve very similar results to FedAvg on these datasets under group-specific distributed concept drift, as they lack the adaptability to effectively address the challenges posed by these evolving data distributions. Nevertheless, Oracle serves as a perfect algorithm for this problem setup and should act as an ideal baseline.
VI Experimental Results and Discussion

### v4

@@ III-C Objective
We present the optimization objective employed to detect and address group-specific distributed concept drift while upholding the privacy of participating clients’ data within the federation. We adopt a multiple global model approach where we aim to discover clusters of clients, each representing a distinct concept, and train each global model accordingly. To achieve this, we incorporate local group-specific loss monitoring on each client. In other words, for each group $s\in S$ we calculate its loss on each global model $\theta_{m}$, where $m$ is the index of each model, in the set of available global models, $GM$, and assign each client to a specific global model, accordingly.
@@
$\begin{split}\min_{\theta_{m}\in GM}f(\theta_{m})^{t}=\frac{\sum_{k\in K}G_{k}(\theta_{m})\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}{\sum_{k\in K}\sum_{t^{\prime}\leq t:w_{k}^{t^{\prime}}=m}|D_{k}^{t^{\prime}}|}\\ w_{k}^{t}=\arg\min_{m}\sum_{s\in S}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \text{s.t.}\quad\forall\theta_{m}\in GM:\forall k:w_{k}^{t}=m:\forall s\in S:\\ \ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{w_{k}^{t-1}})^{D_{k}^{t-1}}_{s}+\delta_{s}\end{split}$ (3)
@@
  • $\theta_{m}\in GM$ refers to a specific global model with index $m$;
  • $w_{k}^{t}$ denotes the index of the global model that client $k$ has identified with at timestep $t$;
@@
  • $f(\theta_{m})^{t}$ is the weighted sum of local objectives $G_{k}(\theta_{m})$ of clients that identify with global model $\theta_{m}$;
@@
  • $\delta_{s}$ represents a predefined threshold for the acceptable loss difference of group $s$ between two consecutive timestamps.
Hence, at each timestep, each client identifies with the model that yields the lowest sum of group losses, as long as the loss difference of each group $s$ between the current and the previous timestamp does not surpass the threshold for that group, $\delta_{s}$. If there is no such model, then a new global model will be created and added to the set of global models $GM$. Further details will be explained in the next section.
Finally, our framework involves training multiple global models, each assigned to a cluster of clients. Within each round, the training of each global model proceeds via standard FedAvg [33] over the clients assigned to it. Since the aggregation weights for clients in a given cluster are fixed during that timestep, the training dynamics are equivalent to those in classical FedAvg. Therefore, existing convergence analysis for FedAvg [34] applies directly within each cluster at each timestep.
@@ IV FairFedDrift: Fair Federated Learning under Group-Specific Distributed Concept Drift
Our work offers a solution to address the challenge of group-specific distributed concept drift building upon an existing algorithm, FedDrift, presented in [20]. Algorithm 1 presents the pseudo-code of FairFedDrift [^1] Source code can be found at: https://github.com/teresalazar13/FairFedDrift..
@@
1: Initialize global models GM: $\left[\theta_{0}\right]$, number of timesteps: $T$, number of rounds: $R$, number of clients: $K$, client identities $w_{k\in K}^{t=0}=0$, local epochs: $E$, local batch size: $|B|$, local learning rate: $\eta$, loss threshold: $\delta_{s}$.
@@
4:   if $\nexists\theta_{m}\in GM:\forall s\in S,\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{w_{k}^{t-1}})^{D_{k}^{t-1}}_{s}+\delta_{s}$ then
@@
   $\sum_{s\in S}\begin{cases}\ell(\theta_{m})^{D_{k}^{t}}_{s}\\ \quad\quad\text{if }\ell(\theta_{m})^{D_{k}^{t}}_{s}\leq\ell(\theta_{w_{k}^{t-1}})^{D_{k}^{t-1}}_{s}+\delta_{s}\\ \infty\quad\text{else}\end{cases}$
@@
[…]
@@
17:   Add new model to GMGM: $\theta_{k}=\frac{\theta_{i}W_{i}+\theta_{j}W_{j}}{W_{i}+W_{j}}$,
@@ IV-1 Multi-model approach
IV-2 Group-specific distributed concept drift detection
@@ IV-4 Training and Model Averaging
To sum up, the choice of considering the loss of each group $s\in S$ in FairFedDrift, in contrast to FedDrift [20], stems from its effectiveness in handling group-specific concept drift as it inherently better suited to capture the nuances of fairness in these scenarios where specific groups may experience concept drift at different timesteps. This is attributed to the fact that, in instances of group-specific concept drift, a global loss metric may not manifest substantial variations. FairFedDrift addresses this gap through three core modifications: (1) during model selection, it uses the sum of group-specific losses rather than global loss to ensure that all groups are considered in the selected model; (2) during drift detection, it monitors each group’s loss individually, flagging drift when any group-specific variation exceeds the threshold; and (3) in model merging, it imposes per-group thresholds to merge models together. These changes are essential for maintaining fairness over time in dynamic, heterogeneous environments. We provide a detailed comparison between FairFedDrift and baseline methods, and empirically validate its advantages in Section VI.
IV-5 Window Size
Finally, instead of retaining the entire history of client identities $\cup_{t^{\prime}\leq t}w_{k}^{t^{\prime}}$ and local datasets $\cup_{t^{\prime}\leq t}D_{k}^{t^{\prime}}$ at each timestamp, clients could opt to maintain only a sliding window of the most recent timesteps [20]. The window size parameter determines how much past information is used during model assignment and concept drift detection. Mathematically, this adjustment limits the history to a window of size $w$, encompassing timestamps in the range $\max(0,t-w)\leq t^{\prime}\leq t$. Specifically, for each timestep $t$, the history of client identities $w_{k}^{t^{\prime}}$ and local datasets $D_{k}^{t^{\prime}}$ are restricted to: $\cup_{\max(0,t-w)\leq t^{\prime}\leq t}w_{k}^{t^{\prime}}$ and $\cup_{\max(0,t-w)\leq t^{\prime}\leq t}D_{k}^{t^{\prime}}$, respectively. This change reduces the computational burden while potentially affecting fairness and accuracy, as less historical data is considered.
@@ V Methodology and Experimental Design
V-A Datasets and Group-Specific Distributed Concept Drift Simulation
In our experimental setup, we leverage diverse datasets derived from established benchmarks in machine learning: MNIST [35], FEMNIST [36], CIFAR-10 [37], and Adult [38]. These datasets were chosen due to their extensive size, making them well-suited for this FL setup, where each client necessitates a substantial number of data instances across different timesteps. In addition, these datasets have also been employed in fairness studies [39, 40], which makes them suitable to our experiments. To simulate this FL scenario, we partition each dataset into $K$ clients and $T$ timesteps and introduce group-specific concept drift within the datasets at specific timesteps with distinct concepts (‘A’, ‘B’, ‘C’, ‘D’, ‘E’).
@@ V-A1 MNIST-GDrift
The MNIST dataset consists of gray-scale handwritten digit images, commonly used for digit recognition tasks [35]. We introduce one sensitive attribute, $S$, with two groups ($S=1$ and $S=0$), representing distinct image characteristics. For $S=0$ images, we invert the background and digit colors and rotate 90 degrees in comparison to standard MNIST images ($S=1$). To simulate four distinctive group-specific concepts (‘B’, ‘C’, ‘D’, ‘E’) from the original distribution (‘A’), we label-swap images for $S=0$: for drift ‘B’, we swap labels ‘1’ and ‘2’; for drift ‘C’, we swap labels ‘2’ and ‘3’; for drift ‘D’, we swap labels ‘3’ and ‘4’; for drift ‘E’, we swap labels ‘4’ and ‘5’.
@@ V-A2 FEMNIST-GDrift
V-A3 CIFAR-10-GDrift
The CIFAR-10 dataset is an established dataset used for object recognition. It is a subset of the 80 million tiny images dataset and consists of color images containing one of 10 object classes. We introduce one sensitive attribute, $S$, with two groups ($S=1$ and $S=0$). Images where $S=0$ correspond to classes 0-4, while images with $S=1$ correspond to classes 5-9. To simulate four distinctive group-specific concepts (‘B’, ‘C’, ‘D’, ‘E’) from the original distribution (‘A’), we label-swap images for $S=0$: for drift ‘B’, we swap labels ‘0’ and ‘1’; for drift ‘C’, we swap labels ‘1’ and ‘2’; for drift ‘D’, we swap labels ‘2’ and ‘3’; for drift ‘E’, we swap labels ‘3’ and ‘4’.
V-A4 Adult-GDrift
@@
We experiment on the Adult dataset [38] commonly used in fair machine learning research. The prediction task for the Adult dataset involves determining whether a person’s income exceeds \$50K/yr. In this context, race serves as the sensitive attribute, where $S=1$ typically corresponds to White individuals, and $S=0$ represents non-White individuals. To introduce concept drift, we modify features of the samples where $S=0$ and change the outcome of each of 4 distinct groups created based on gender (‘Male’ and ‘Female’) and occupation (‘Professional’ - ‘Adm-clerical’, ‘Exec-managerial’, ‘Prof-specialty’, ‘Tech-support’; and ‘Technical’ - ’Armed-Forces’, ‘Craft-repair’, ‘Machine-op-inspct’, ‘Farming-fishing’, ‘Handlers-cleaners’, ‘Other-service’, ‘Priv-house-serv’, ‘Protective-serv’, ‘Sales’, ‘Transport-moving’). For concept ‘A’, the dataset remains unchanged. For concept ‘B’, the target class of non-White individuals who are Male and have ‘Professional’ occupations is set to 1. For concept ‘C’, the target class of non-White individuals who are Female and have ‘Professional’ occupations is set to 1. For concept ‘D’, the target class of non-White individuals who are Male and have ‘Technical’ occupations is set to 1. For concept ‘E’, the target class of non-White individuals who are Female and have ‘Technical’ occupations is set to 1.
V-A5 $\alpha$ parameter
@@
The parameter $\alpha$ is an important element in our experimental design, determining the size of the unprivileged group relative to the privileged group. The size of the unprivileged group is a result of multiplying the size of the privileged group by $\alpha$. By adjusting the value of $\alpha$, we can control the balance between the groups in our datasets, allowing us to explore different scenarios where one group may be underrepresented or over-represented relative to the other. As such, when $\alpha$ is closer to zero, the relative size of the unprivileged group ($S=0$) is smaller. To generate a specific number of instances depending on the value of $\alpha$ for the MNIST-GDrift and FEMNIST-GDrift datasets, we adjust the number of transformed images accordingly. For the CIFAR-10-GDrift dataset, we reduce the number of images from the unprivileged group ($S=0$, classes 0-4) by sampling the dataset based on the specified $\alpha$ value, while keeping all images for the privileged group ($S=1$, classes 5-9). Finally, for the Adult-GDrift dataset, the process is more intricate as the original dataset is oversampled, whose default representation is equivalent to $\alpha=0.16$. If $\alpha>0.16$, we generate unprivileged instances, and if $\alpha<0.16$, we generate privileged instances. We employ the SMOTE [41] algorithm for creating these instances.
V-A6 Scenarios
We design five distinct scenarios of group-specific distributed concept drift with distinct concepts (‘A’, ‘B’, ‘C’, ‘D’, ‘E’), as illustrated in Figure 2. These aim to provide a comprehensive exploration of the temporal and spatial dynamics that can impact fairness and accuracy in FL settings. More specifically, each scenario is constructed to vary in complexity based on three key factors:
  • Number of distinct concepts – The number of different concepts (i.e. ‘A’, ‘B’, ‘C’, ‘D’ and ‘E’) appearing in the scenario.
  • Number of concept drifts – The number of times a concept drift occurs across all clients.
  • Number of timesteps with drift – The number of timesteps where at least one client experiences a drift.
Below, we provide a more explicit distinction between the scenarios:
  • Scenario 4.1 – Involves three distinct concepts appearing sequentially, with 15 drift events across three timesteps.
  • Scenario 4.2 – Also has three concepts, but more recurring drifts occur, leading to 18 drift events across the same three timesteps.
  • Scenario 4.3 – Introduces simultaneous concept appearances, increasing complexity to 23 drift events across four timesteps.
  • Scenario 4.4 – Expands to five distinct concepts, resulting in 30 drift events across six timesteps, with more diverse transitions.
  • Scenario 4.5 – The most complex setting, featuring five concepts and a highly unpredictable pattern, with 38 drift events across seven timesteps.
By evaluating performance across this spectrum of drift scenarios, we ensure that our approach is tested under both controlled and highly dynamic drift conditions, reflecting the diverse challenges found in real-world FL applications.
@@ V-B Implementation Details
We employ a convolutional neural network (CNN) for the MNIST-GDrift and FEMNIST-GDrift datasets featuring a convolutional layer, a max-pooling layer, a fully connected layer, and a softmax activation function. For the CIFAR-10-GDrift dataset we use a ShuffleNet v2 [42] model pre-trained on ImageNet, and extend it by adding a fully connected layer with 256 units, followed by ReLU, dropout, batch normalization, another fully connected layer, using softmax for multi-class classification. For the Adult-GDrift dataset we employ a fully connected feed-forward neural network with one hidden layer with 10 tanhtanh units, and a sigmoid output neuron (binary classification). As explained before, each test scenario is constructed to span $T=10$ timesteps, with $K=10$ clients participating. At each timestep, clients perform $R=10$ local communication rounds using a learning rate of $\eta=0.1$. We experiment with different values of $\delta_{s}$: $\{0.05,0.1,0.25,0.5,0.75,1.0,1.25,1.5,1.75,2.0\}$. Furthermore, for the MNIST-GDrift and FEMNIST-GDrift datasets we set $|B|=32$ and $E=5$. For the CIFAR-10-GDrift dataset we set $|B|=128$ and $E=30$. For the Adult-GDrift dataset we set $|B|=10$ and $E=10$.
@@ V-C Metrics
We evaluate the algorithms based on both accuracy (ACC) and fairness (AEQ, OEQ, OPP).
@@
Accuracy Equality (AEQ) [43] is a measure of fairness that calculates the ratio of accuracies across two groups ($S=0$ and $S=1$):
@@
Overall Equality of Opportunity (OEQ) [44] compares the true positive rates (TPRs) for two groups. Let $TPR_{0}=P[\hat{Y}=c\mid S=0,Y=c]$ and $TPR_{1}=P[\hat{Y}=c\mid S=1,Y=c]$ represent the TPRs for group $S=0$ and $S=1$ for the $c^{th}$ target class, respectively. The formula for OEQ is given by:
$\text{OEQ}=\left\{\begin{array}[]{ll}\frac{1}{C}\sum_{c=1}^{C}\frac{TPR_{0}}{TPR_{1}}&\text{If}\quad Y(S=0)=Y(S=1)\vskip 8.5359pt\\ \frac{\frac{1}{C_{0}}\sum_{c=i}^{C_{0}}TPR_{0}}{\frac{1}{C_{1}}\sum_{c=j}^{C_{1}}TPR_{1}}&\text{If}\quad Y(S=0)\neq Y(S=1)\vskip 8.5359pt\end{array}\right$. (5)
where $C$ is the number of target classes. In cases where the target classes for both groups are overlapping (i.e. $Y(S=0)=Y(S=1)$), the metric compares the TPRs across the groups for each class. However, if the classes are non-overlapping (i.e. $Y(S=0)\neq Y(S=1)$), as in the CIFAR-10-GDrift dataset, the metric evaluates the TPRs for each class separately within each group and then computes the ratio of the average TPRs across groups. Here, $C_{0}$ is the number of classes $c$ where the count of instances with $S=0$ and $Y=c$ is nonzero (where $TPR_{0}$ can be calculated), and $C_{1}$ is defined similarly for $S=1$.
@@
Overall Predictive Parity (OPP) [45] compares the positive predictive values (PPVs) for two groups. Let $PPV_{0}=P[Y=c\mid S=0,\hat{Y}=c]$ and $PPV_{1}=P[Y=c\mid S=1,\hat{Y}=c]$ represent the PPVs for group $S=0$ and $S=1$ for the $c^{th}$ target class, respectively. The formula for OPP is given by:
$\text{OPP}=\left\{\begin{array}[]{ll}\frac{1}{C}\sum_{c=1}^{C}\frac{PPV_{0}}{PPV_{1}}&\text{If}\quad Y(S=0)=Y(S=1)\vskip 8.5359pt\\ \frac{\frac{1}{C_{0}}\sum_{c=i}^{C_{0}}PPV_{0}}{\frac{1}{C_{1}}\sum_{c=j}^{C_{1}}PPV_{1}}&\text{If}\quad Y(S=0)\neq Y(S=1)\vskip 8.5359pt\end{array}\right$. (6)
where $C$ is the number of target classes. In a similar way to the OEQ metric, $C_{0}$ is the number of classes $c$ where the count of instances with $S=0$ and $\hat{Y}=c$ is nonzero (where $PPV_{0}$ can be calculated), and $C_{1}$ is defined similarly for $S=1$.
@@
V-D Baselines
FairFedDrift is compared against three pivotal algorithms: FedAvg, FedDrift [20] and Oracle. The Oracle algorithm acts as an idealized benchmark, assuming access to concept IDs during training and using multiple-model training based on ground-truth clustering. While unrealistic in real-world scenarios, where concept IDs are unknown, Oracle provides an upper-bound performance reference to measure the effectiveness of practical approaches [20]. Given the unique problem of group-specific distributed concept drift, traditional baselines for fair FL are absent, as explained in Section II. We found that these achieve very similar results to FedAvg on these datasets under group-specific distributed concept drift, as they lack the adaptability to effectively address the challenges posed by these evolving data distributions. Nevertheless, Oracle serves as a perfect algorithm for this problem setup and should act as an ideal baseline.
VI Experimental Results